Calculus I, Fall 2012 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 13

1. Section 6.11: Problems: 17, 21, 23, 27.

2. Section 7.1: Problems: 9, 37, 47.

3. Section 7.2: Problems: 3, 5, 41.

4. Section 7.3: Problems: 43, 45, 46.